OpenStudy (anonymous):

you are building a rectangular dog pen with an area of 90 ft^2. you want the length of the pen to be 3 ft longer than twice its width. which equation can you use to find the width of the pen?

OpenStudy (anonymous):

Area = length * width = 90 but you want "length of the pen to be 3 ft longer than twice its width" so you need to write the length in terms of the width... so, length = 2*width + 3 so area is now: length * width = 90 (2*width + 3) * width = 90 (2w + 3)w = 90 can you solve this equation from here?